In a particular city, one in three families have a phone in their home. If 90 families are chosen at random, calculate the probability that at least 30 of them will have a phone.


1
Expert's answer
2021-10-08T08:06:48-0400

Let "X=" the number of families which have a phone: "X\sim Bin(n, p)."

Given "n=90, p=\dfrac{1}{3}, q=1-p=1-\dfrac{1}{3}=\dfrac{2}{3}."


"P(X\geq30)=1-P(X<30)"

"=1-\displaystyle\sum_{x=0}^{29}\dbinom{90}{x}(\dfrac{1}{3})^{x}(\dfrac{2}{3})^{90-x}"

"\approx0.53956753215""np=90(\dfrac{1}{3})=30\geq10, nq=90(\dfrac{2}{3})=60\geq10"

Then "X" has approximately a normal distribution with "\mu=np=30" and "\sigma=\sqrt{npq}=\sqrt{90(\dfrac{1}{3})(\dfrac{2}{3})}=2\sqrt{5}."


"P(X\geq30)=P(X>29.5)=1-P(X\leq29.5)"

"=1-P(Z\leq\dfrac{29.5-30}{2\sqrt{5}})\approx1-P(Z\leq-0.1118)"

"\approx0.544509"